answer to (9,3),(-6,7y);m=3 using coordinate slope of a line

Answer by mananth(16949)

(9,3),(-6,7y);m=3 using coordinate slope of a line
...
slope m = y2-y1 / x2-x1
7y-3 / -6-9 = m
7y-3 / -15 = 3
7y-3=-45
7y=-45+3
7y=-42
/7
y=-6
...
m.ananth@hotmail.ca